OpenStudy (anonymous):

Jackie scored 82 and 88 on her first two quizzes. Write and solve a compound inequality to find the possible values for a third quiz score that would give her an average between 80 and 85, inclusive.

OpenStudy (anonymous):

\[A. 80 \le \frac{ 82+88+n }{ 3 }\le 85; 70\le n \le 85\]

OpenStudy (anonymous):

b.\[B. 70\le \frac{ 82+88+n }{ 3 }\le 85; 80\le n \le 85\]

OpenStudy (anonymous):

\[C. 80< \frac{ 82+88+n }{ 3 }<85; 70 <n<85\]

OpenStudy (anonymous):

\[D. 82\le \frac{ 80+85+n }{ 3 }\le 88; 70len \le85\]

OpenStudy (anonymous):

^theses are the answers A,B,C,and D

OpenStudy (anonymous):

Since its multiple choice, you could just plug in n to the fraction and see if you get 80-85. For D its incorrect because it added 80 and 85 instead of 82 and 88. Now just look at the inequality symbols, you want to be between 80 and 85 so\[80\le \frac{ 82+88+n }{ 3 }\le85\]Now, you can eliminate B. It says you want to have a total score between 80 and 85 which means including '80' and '85'. So you finally eliminate C as well. So the answer should be A. If I made a mistake, sorry it's been awhile.

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

\[-\frac{ 2 }{ 5 }x - 9 < \frac{ 9 }{ 10 }\]